Please I would appreciate anyone's help. I am stuck on this problem.
Solve by elimination or substitution. For all three variables a, b and c. Use all three equations.
5a-5b+5c=25
-4a+4b-2c=-12
-a+ b-5c=-21

Answer by ewatrrr(24785)

5a-5b+5c=25
-a+ b-5c=-21
4a - 4b = 4
.........5a-5b+5c=25 0r a-b+c= 5 ***
2(a-b+c= 5)
-4a+4b-2c=-12
-2a + 2b = -2
.......
4a - 4b = 4
2(-2a + 2b = -2)
0 = 0
Infinite Solutions: 4a - 4b = 4 0r a - b = 1
Re TY: Note: Using b = a-1
b = a-1, c = b-a+5 0r c = (a-1) - a + 5 = 4
(a, a-1, 4)